OpenStudy (anonymous):

Part 1: Solve the inequality -5x + 8 less than or greater to x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol.

OpenStudy (anonymous):

less than or greater to? Is that supposed to say less than or equal to?

OpenStudy (anonymous):

\[-5x + 8 \le x - 4\] subtract x from each side \[-6x + 8 \le -4\] subtract 8 from each side \[-6x \le -12\] divide by -6 \[x \le 2\] there ya go

OpenStudy (anonymous):

greater too. sorry (:

OpenStudy (anonymous):

Ashley, based on my explanation to you in the other thread, see if you can find Scramble's mistake.

OpenStudy (anonymous):

I cant find it 0.0

OpenStudy (anonymous):

You can solve an inequality for x the same way you would solve an equation for x with only 1 little difference. If you multiply or divide both sides of the equation by a negative, then you have to switch the direction of the inequality arrow.

OpenStudy (anonymous):

Ohh. thank you! I get it now(:

OpenStudy (anonymous):

=D You're welcome!

OpenStudy (anonymous):

=)